OpenStudy (thecatman):

A hotel lobby can seat 14 people at chairs and loveseats. The lobby has a total of 11 pieces of furniture for seating. Assuming a chair seats 1 person and a loveseat seats 2, how many loveseats are in the lobby? 2 3 4 5

OpenStudy (paki):

Let L be the # of loveseats 14 = 2*(L) + 1(11 - L) 14 = 2L + 11 - 1L 14 - 11 = 2L - 1L 3 = 1L or L = 3... @TheCatMan